OpenStudy (anonymous):

Help now please! What is the sum of 9/26 and -4/13? Thank you so much!

OpenStudy (anonymous):

You need to form a least common denominator.

OpenStudy (anonymous):

26?

OpenStudy (anonymous):

is the answer one?

OpenStudy (anonymous):

9/26 - 8/26 =1?

OpenStudy (anonymous):

Oh so close. 9/26 - 8/26 = 1/26

OpenStudy (anonymous):

Good work

OpenStudy (anonymous):